OpenStudy (anonymous):

if the perimeter of a rectangle is p and its diagonal is d, the difference between the lenght and width of the rectangle is?

OpenStudy (pfenn1):

Let x and y equal the lengths of the two sides of the rectangle. What is the formula for the perimeter p based on x and y . If the diameter is d, what would be the formula for that in terms of x and y.

OpenStudy (anonymous):

the formula????

Parth (parthkohli):

 The formula for perimeter is 2(l + b) So, basically, 2(l + b) can be said as the perimeter. Now, you'd know the Pythagorean theorem. It says that a^2 + b^2 = c^2 We can say that d^2 = l^2 + b^2 Continue now.

OpenStudy (pfenn1):

Perimeter p = 2x + 2y. The perimeter is the sum of the lengths of the 4 sides.

OpenStudy (anonymous):

i am confused so if the formula is p = 2x + 2y, then what do i do?

OpenStudy (anonymous):

As ParthKohli has mentioned, \[d ^{2} = l ^{2} + b ^{2}\] and \[p = 2l + 2b\] Can you find l - b now?

OpenStudy (anonymous):

yeay thats much simpler thanks!

OpenStudy (anonymous):

Actually, even I'm confused now, how do you proceed?

OpenStudy (anonymous):

p = 2(l + b) = -2(-l - b) -l - b = -p/2

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

oh wait i thought you use the Pythagorean theorem to solve the rest??

OpenStudy (anonymous):

this problem is hopeless. i'm going to kill my teacher

OpenStudy (anonymous):

I'm not so sure ... why don't we take the advice of @ParthKohli?

OpenStudy (apoorvk):

No please don't kill your teacher. lol are you familiar with pythagoras theorem?

OpenStudy (anonymous):

do you mean the Pythagorean theorem???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

EVERYONE knows Pythagorean theorem .

OpenStudy (pfenn1):

Did your teacher give you any other information like what d and p are?

OpenStudy (anonymous):

sadly no,

OpenStudy (pfenn1):

Wow, I have been trying this all different ways and haven't gotten anywhere with it. Maybe we should ask @FoolForMath to step in and help.

OpenStudy (anonymous):

i am giving up. thanks for trying help out you guys

OpenStudy (pfenn1):

Sorry. That one is a stumper.

OpenStudy (pfenn1):

@Zarkon, can you help?

Parth (parthkohli):

Lol I'm not sure what I have to do here

OpenStudy (pfenn1):

@ParthKohli We can't seem to get this, @tim538 seems to have given up but it bugs me that I can't get it.

Parth (parthkohli):

Maybe the difference of the lengths and width is l - b lolz

OpenStudy (anonymous):

i'll show you link, it number 4 i don't know if this helps but you can see original problem for yourself.

Parth (parthkohli):

I'm pretty sure about that one

OpenStudy (anonymous):

2(l+b)=p => (l+b)=p/2 => l^2 +b^2+2lb= p^2/4 => 2lb= p^2/4-d^2 (l-b)^2= l^2+b^2-2lb => \( (l-b)= \sqrt{2d^2-p^2/4} \)

OpenStudy (anonymous):

i said 3 instead of 4

OpenStudy (anonymous):

@FoolForMath you got it that easily!!!

OpenStudy (anonymous):

;)

OpenStudy (pfenn1):

From the formula for the perimeter\[2(x+y)=p\]Rearrange and square both sides\[(x+y)=p/2\]\[(x+y)^2=\frac {p^2}{2^2}\]\[x^2 +y^2+2xy= p^2/4\]Rearrange again to solve for 2xy\[2xy=\frac {p^2}{4}-(x^2+y^2)\]Remember and substitute \[x^2 + y^2=d^2\]\[2xy=\frac{p^2}{4}-d^2\]\[(x-y)^2= x^2+y^2-2xy\]\[(x-y)^2=d^2-\left(\frac{ p^2}{4}-d^2 \right)=-\frac {p^2}{4}\]\[(x-y)= \sqrt{-p^2/4}\]

OpenStudy (pfenn1):

@FoolForMath, does the d^2 term drop at the end?

OpenStudy (pfenn1):

*doesn't

OpenStudy (inkyvoyd):

I don't know the pythagorean theorem. Trolol

OpenStudy (anonymous):

\[(x-y)^2=d^2-\left(\frac{ p^2}{4}-d^2 \right)=d^2-\frac {p^2} 4+d^2 =2d^2 -\frac {p^2}{4} \]

OpenStudy (pfenn1):

@FoolForMath, yes, math mistake. Thanks for clearing that up.